An investment project provides cash inflows of $770 per year for 11 years.  

a. What is the project payback period if the initial cost is $3,080?

b. What is the project payback period if the initial cost is $4,543?

c. What is the project payback period if the initial cost is $9,240?
Answer #1

Payback period=initial cost/annual cash flows

1.Payback period=(3080/770)=4 years

2.Payback period=(4543/770)=5.9 years

3.Payback period=(9240/770)=0